# An Empirical Study of Testing Practices in Open Source AI Agent Frameworks and Agentic Applications

### v2

@@ Abstract
To address this gap, we conduct the first large-scale empirical study of testing practices in the AI agent ecosystem, analyzing 39 open-source agent frameworks and 439 agentic applications. We identify ten distinct testing patterns and find that novel, agent-specific methods like DeepEval are seldom used (around 1%), while traditional patterns like negative and membership testing are widely adapted to manage FM uncertainty. By mapping these patterns to canonical architectural components of agent frameworks and agentic applications, we uncover a fundamental inversion of testing effort: deterministic components like Resource Artifacts (tools) and Coordination Artifacts (workflows) consume over 70% of testing effort, while the FM-based Plan Body receives less than 5%. Crucially, this reveals a critical blind spot, as the Trigger component (prompts) remains neglected, appearing in around 1% of all tests.
@@ 1 Introduction
With the proliferation of agent frameworks and agentic applications, along with the increasingly critical tasks agents are asked to perform, evaluating their effectiveness has become essential. The dominant evaluation culture, inherited from traditional AI, relies on standardized benchmarks that measure task performance (Liu et al., 2023a). For example, AgentBench introduced a multi-dimensional benchmark with eight distinct interactive environments to systematically assess an LLM-as-agent’s reasoning and decision-making abilities (Liu et al., 2023a). A similar approach is followed by other popular benchmarks, e.g., GAIA (Mialon et al., 2023), WebArena (Zhou et al., 2023), and MINT (Wang et al., 2023). While valuable for objective comparison, this approach creates a dangerous gap between perceived performance and real-world reliability. Benchmarks primarily test if an agent can succeed on a set of pre-defined tasks, but not whether it is robust, safe, or dependable for an end user’s specific needs. For instance, an agent that tops a leaderboard might still perform poorly when faced with edge cases, get stuck in buggy loops, produce hallucinations, or succumb to unhandled faults that benchmarks rarely account for (Weidinger et al., 2025).
To address this reliability gap, practitioners increasingly adopt traditional software testing methods, e.g., unit testing, to verify agent behavior in controlled scenarios. Unlike benchmarks, unit tests not only verify that the software behaves as intended, but also expose edge cases and uncover newly introduced faults in code chunks that previously worked properly (Niedermayr et al., 2016). However, adapting traditional testing practices to FM-based agent frameworks and agentic applications is tricky due to the inherent non-determinism and non-reproducibility of the underlying models (Hassan et al., 2024).
@@
Surprisingly, this critical area remains almost entirely unexplored. While research has focused extensively on agent architectures, capabilities, and benchmarks, there is a clear lack of empirical evidence on how practitioners actually test agent frameworks and agentic applications. We found no comprehensive studies examining the state-of-the-art testing practices for open-source agents: how developers structure tests, verify non-deterministic outputs, and which parts of the agent they prioritize while testing. These gaps are concerning, given that the potential impact of untested or unverified agents can be poor in real-world use cases. For example, previous studies have already reported that the performance of FM-based systems can deteriorate during model upgrades, and without testing the prompts and managing the versions, it is difficult to identify such silent performance degradation (Ma et al., 2024). Similarly, prior studies have reported at least seven error patterns while invoking tools in FM-based systems (Kokane et al., 2025), which can also lead to production issues if not appropriately tested during the development phase. To complement the findings from researchers, we have also observed that different components (e.g., prompts, tools) behave differently in different FMs or even on different versions of the same FM. [^1] https://community.openai.com/t/gpt-4o-vs-gpt-4-turbo-function-calling/750493 and https://www.reddit.com/r/ChatGPT/comments/1idghel/was_anyone_elses_experience_with_gpt4o_completely
@@
The challenge is further amplified by the rapid evolution of these agent frameworks and agentic applications. Practitioners are confronted with a constant flux of new, implementation-specific components, e.g., from novel tool invocation protocols and multi-agent communication standards (MCP, A2A, ACP) to evolving memory and planning modules (Hasan et al., 2025; Ehtesham et al., 2025). This architectural evolution makes it exceedingly difficult to establish durable testing strategies; what is tested today may become obsolete or change fundamentally tomorrow. As established in software engineering research, the key to managing such complexity is to map these volatile, concrete components to a stable, canonical or conceptual architectural model, which provides a durable foundation for reasoning about system quality and test coverage (Lukyanenko et al., 2024). While prior work has proposed valuable taxonomies that categorize emerging agent components (Händler, 2023a), we have not seen any work grounding the architecture component to any conceptual architecture till date.
@@
To fill these critical gaps, we conduct the first large-scale empirical study of unit testing practices in the agent ecosystem. We analyze 39 agent frameworks and 439 agentic applications, examining how practitioners arrange tests and assert correctness. Then we map the state-of-the-art agent architecture component to an existing conceptual architecture framework Boissier et al. (2020) and identify component-specific testing patterns. Our study answers the following research questions (RQs):
@@
Motivation: The growing adoption of FM-based agents in business-critical domains has intensified the need for robust quality assurance. However, the inherent non-determinism and non-reproducibility of foundation models (FMs) significantly complicate the application of traditional, deterministic testing techniques (Hassan et al., 2024). These challenges require practitioners to adopt alternative testing strategies that can accommodate unpredictable behavior during both test arrangement and result verification.
Despite extensive research into testing methodologies within traditional software engineering (Zhu et al., 2025; Zamprogno et al., 2022) and machine learning applications (Openja et al., 2024), the FM-based agent ecosystem remains underexplored. This study fills that gap by conducting the first large-scale empirical investigation of unit testing practices in open-source agent frameworks and applications. By systematically analyzing how practitioners structure and verify test functions under these constraints, the work provides a foundational catalog of testing strategies that can guide the development of more reliable and resilient agentic systems.
@@
Motivation: Agent ecosystem is a complex ecosystem of interacting components, from memory and tools to planning and coordination artifacts (Hettiarachchi, 2025). Each of these components introduces distinct behaviors and potential failure modes, demanding tailored testing strategies. Understanding how testing effort is distributed across this heterogeneous architecture is vital for assessing whether critical components receive sufficient attention and for supporting the development of more systematic and effective testing practices. A comprehensive empirical understanding of component-wise testing patterns is also essential for helping practitioners formulate robust testing strategies for new agent systems.
A significant challenge for empirical study, however, is the ecosystem’s rapid evolution, with standards for core functions like tool use (MCP) Hasan et al. (2025), communication (A2A, ACP, ANP) (Ehtesham et al., 2025), and memory emerging and changing continuously. To create a durable analysis that transcends these rapid implementations, we opt to map observed components to a stable, canonical architectural model, e.g., JaCaMo framework Boissier et al. (2020), which will allow us to assess whether testing effort is distributed effectively across fundamental roles, supporting the development of more systematic testing practices. To this end, our second research question first maps the observed components in the test functions to canonical components of conceptual agent architecture and then investigates the relationship between the testing patterns identified in RQ1 and the underlying components of agent frameworks and applications.
@@ 2 A Motivational Example
John, a senior developer with a background in traditional software engineering, recently joined an ambitious AI startup. His first assignment is to lead the development of an agentic application: a playful, interactive voice assistant for children. The agent’s core function is to entertain by dynamically generating jokes and bedtime stories using a state-of-the-art Foundation Model (FM).
Scenario 1: It Just Works. Using a popular open-source agent framework, John quickly builds a prototype. The framework provides seamless integration with the FM. When a child asks for a story, the agent prompts the FM, which returns a creative and engaging narrative. The initial results are impressive, and the team moves forward with deployment, confident in the power of their new AI-driven architecture.
Scenario 2: The Unseen Decay. Weeks after launch, user feedback reveals a degradation in quality. Jokes are becoming boring, and stories occasionally lack coherence. John investigates and finds the root cause: a silent, unannounced update to the underlying FM has altered its interpretation of their prompts. The application is performing, yet the user-facing quality has declined. He realizes he has no automated way to detect this semantic drift, e.g., no tests. He wonders: How do other teams test for this?
Scenario 3: The Component Stack-Up. Things become more complicated with the product team introducing two new features:
@@
    Contextual Follow-up: Children should be able to ask for a story “just like the one from last night.” This feature requires integrating a VectorDB for Retrieval-Augmented Generation (RAG) to maintain conversational history and ensure story similarity (Lewis et al., 2020).
@@
    Real-World Q&A: During a story about Cinderella, a child might ask, “How much is a ticket to Disneyland?” This requires the agent to use a third-party search tool, e.g., an MCP server (Hasan et al., 2025).
John’s testing challenges have now multiplied. Not only does he need to validate the non-deterministic output of the FM, but he must now also evaluate distinct components:
@@
    The RAG system: How can he write a test to verify that a newly generated story is “similar and relevant” to a previous one? What metrics define “similarity”?
@@
    The MCP Server: How does he test a third-party MCP Server for accuracy, fairness, and truthfulness? The server is a black box that could change its behavior at any time.
Scenario 4: The Search for a Standard. Coming from traditional software engineering, John is familiar with mocking, stubbing, and integration testing. He attempts to apply these concepts, but they feel ill-suited for a system with many non-deterministic and independently evolving parts. He investigates how other, more mature open-source agentic applications and frameworks handle this. He dives into the GitHub repositories of popular projects, looking for their testing suites, hoping to find established best practices or community-standard tools for evaluating these complex components.
Scenario 5: The Challenge. John’s search leaves him perplexed. He finds a scattered landscape of ad-hoc scripts, manual evaluation checklists, and a heavy reliance on costly, slow, end-to-end testing that offers little insight into which component failed. There appears to be no standardized methodology for ensuring the quality and reliability of agentic applications. This leads him to a critical set of questions that must be answered to mature the field of agent development:
@@
    How are the components of AI agent frameworks and agentic applications evaluated?(explored in RQ2)
@@ 3.1 Foundations of Software Testing
3.1.1 Testing
In software engineering, software testing is a core part of quality assurance. It is empirically defined as the systematic execution of a program or system with the explicit objective of identifying and isolating defects (Myers, 2006). The fundamental purpose of this process is to methodically evaluate the functional and non-functional attributes of the software, thereby ensuring its adherence to specified requirements prior to its release and implementation.
3.1.2 Subject Under Test (SUT)
The Subject Under Test (SUT) is the specific software artifact, e.g., a function, class, module, or entire service, that is being validated in the testing process (Meszaros, 2007). In our study of the agent ecosystem, SUTs range from low-level resource handlers and parsers to complex, high-level agentic workflows.
3.1.3 Test Function
A test function, also known as a test case, is the atomic unit of verification. It is an executable procedure designed to validate a specific behavior or functionality of the SUT (Van Rompaey and Demeyer, 2008). In practice, a test function initializes a specific state or fixture, executes a command on the SUT, and verifies the outcome using one or more assertions (Tao, 2009).
3.1.4 Arrange-Act-Assert (AAA)
The Arrange-Act-Assert (AAA) pattern is a widely adopted convention for structuring test functions to enhance their readability and maintainability (Wei et al., 2022a). The pattern divides a test into three distinct phases: arranging the preconditions and inputs, acting on the SUT, and asserting that the outcome meets expectations. We adopt AAA as the conceptual framework for analyzing and classifying the testing practices in our study.
3.1.5 Testing Patterns
Testing patterns are recurring, expert solutions to common problems in test design and implementation (Meszaros, 2007). These patterns provide proven strategies for tasks such as isolating the SUT from its dependencies, managing test data, and verifying complex outcomes. For instance, the use of test doubles (e.g., mocks or stubs) is a classic pattern for isolating a SUT (Zhu et al., 2025). These patterns improve the manageability and reliability of a test suite.
@@ 3.2 The Agent Ecosystem
3.2.1 Agent
An agent is traditionally defined as an autonomous entity that executes decisions and actions within a given environment to realize predetermined objectives (Liu et al., 2023a). This conceptualization has its roots in established paradigms such as the Belief-Desire-Intention (BDI) framework (Rao et al., 1995). The capabilities of such agents have been significantly amplified by the development of Foundation Models (FMs). In contemporary implementations, agents leverage a core FM that functions as a cognitive engine, which is supplemented by essential modules for memory, strategic planning, and the utilization of external tools. This integration empowers the agents to address intricate and unstructured challenges with a high degree of autonomy (Park et al., 2023; Hettiarachchi, 2025).
3.2.2 Agent Framework
Agent frameworks are software libraries that provide a structured environment for building, orchestrating, and deploying agentic systems. Frameworks like AutoGen (Wu et al., 2024a) and CAMEL (Li et al., 2023) abstract away the underlying complexities of state management, tool integration, and multi-agent communication. By offering high-level APIs for defining agent roles and workflows, they accelerate the development of sophisticated single and multi-agent applications (Händler, 2023b).
3.2.3 Agentic Application
In this study, an agentic application is a software system built on top of agent frameworks. These applications leverage FMs to perform complex tasks that often require interaction with external systems, such as a storyteller agent that autonomously invokes APIs for searching characters, booking tickets, and querying vectorDBs as shown in the motivational example of Section 2. Such applications use the capabilities provided by the agent frameworks described above to manage their operational logic.
@@ 3.3 Canonical Agent Architecture
3.3.1 Belief Base
3.3.2 Goal
3.3.3 Plan
3.3.4 Trigger
3.3.5 Context
3.3.6 Plan Body
3.3.7 Internal Action
3.3.8 External Action
3.3.9 Communicative Action
3.3.10 Resource Artifact
3.3.11 Coordination Artifact
3.3.12 Boundary Artifact
3.3.13 Observable Property & Signal
3.3.14 Registry

### v3

@@ Abstract
To address this gap, we conduct the first large-scale empirical study of testing practices in the AI agent ecosystem, analyzing 39 open-source agent frameworks and 439 agentic applications. We identify ten distinct testing patterns and find that novel, non-determinism-specific methods like DeepEval are seldom used (around 1%), while traditional patterns like negative and membership testing are widely adapted to manage FM uncertainty. By mapping these patterns to canonical architectural components of agent frameworks and agentic applications, we uncover a fundamental inversion of testing effort: deterministic components like Resource Artifacts (tools) and Coordination Artifacts (workflows) consume over 70% of testing effort, while the FM-based Plan Body receives less than 5%. Crucially, this reveals a critical blind spot, as the Trigger component (prompts) remains neglected, appearing in around 1% of all tests.
@@ 1 Introduction
With the rapid proliferation of agent frameworks and agentic applications, and the increasingly consequential tasks being delegated to them, evaluation has become a central organizing practice in the community. Most prior work adopts a benchmark-driven evaluation paradigm that measures agent performance on standardized task suites, enabling reproducible comparison under fixed conditions, e.g., AgentBench, GAIA, WebArena, and MINT (Liu et al., 2023a; Mialon et al., 2023; Zhou et al., 2023; Wang et al., 2023). While such benchmarks are effective for objective comparison and assessing task-level capability, they are an incomplete proxy for real-world reliability. In practice, benchmark success does not guarantee robustness to distribution shifts, resilience to tool or environment failures, or stable behavior under underspecified or adversarial user inputs. Consequently, agents that rank highly on leaderboards may still perform poorly when faced with edge cases, get stuck in buggy loops, produce hallucinations, or succumb to unhandled faults when presented with minor variations unseen in the benchmarks (Gevers et al., 2025).
This gap highlights the need to distinguish evaluation from testing in the context of FM-based agents. While often used interchangeably, testing and evaluation diverge significantly in the context of FM-based agents. Evaluation of FM-based agents assesses the interaction settings, agent behavior, and performance, including contextual consistency, factual accuracy, bias, and safety (Ma et al., 2025). In contrast, software testing is an engineering process aimed at verifying that an implemented system behaves as intended, continues to satisfy specified requirements as the code evolves, and exposes newly introduced faults in code chunks that previously worked properly (Niedermayr et al., 2016; Mohammadi et al., 2025). Despite the central role of testing in mature software systems, there remains a lack of systematic study of how practitioners test FM-based agent frameworks and agentic applications in practice.
@@
While research has focused extensively on agent architectures, capabilities, and benchmarks, there is a clear lack of empirical evidence on how practitioners actually test agent frameworks and agentic applications. We found no comprehensive studies examining the state-of-the-art testing practices for open-source agents: how developers structure tests, verify non-deterministic outputs, and which parts of the agent they prioritize while testing. These gaps are concerning, given that the potential impact of untested or unverified agents can be poor in real-world use cases. For example, previous studies have already reported that the performance of FM-based systems can deteriorate during model upgrades. Moreover, we have also observed that different components (e.g., prompts, tools) behave differently in different FMs or even on different versions of the same FM (Ma et al., 2024). Even with the same version of FM, a slight change in prompt phrasing or order can introduce significant behavioral drift (Razavi et al., 2025; Zhuo et al., 2024; Rafi et al., 2026). Without testing, it is difficult to identify such silent performance degradation. Similarly, prior studies have reported at least seven error patterns while invoking tools in FM-based systems (Kokane et al., 2025), which can also lead to production issues if not appropriately tested during the development phase.
@@
The rapid evolution of these agent frameworks and agentic applications further amplifies the challenge. Practitioners are confronted with a constant flux of new, implementation-specific components, e.g., from novel tool invocation protocols and multi-agent communication standards (MCP, A2A, ACP) to evolving memory and planning modules (Hasan et al., 2025; Huang and Hughes, 2025). This architectural evolution makes it exceedingly difficult to establish durable testing strategies; what is tested today may become obsolete or change fundamentally tomorrow. As established in software engineering research, the key to managing such complexity is to map these volatile, concrete components to a stable, canonical, or conceptual architectural model, which provides a durable foundation for reasoning about system quality and test coverage (Lukyanenko et al., 2024). While prior work has proposed valuable taxonomies that categorize emerging agent components (Händler, 2023), we have not seen any work grounding the architecture component to any conceptual architecture till date.
@@
To fill these critical gaps, we conduct the first large-scale empirical study of unit testing practices in the agent ecosystem. We analyze 39 agent frameworks and 439 agentic applications, examining how practitioners arrange tests and assert correctness. Then we map the state-of-the-art agent architecture component to an existing conceptual architecture framework (Boissier et al., 2020) and identify component-specific testing patterns. Our study answers the following research questions (RQs):
@@
Motivation: Adapting traditional testing practices to FM-based agent frameworks and agentic applications is challenging due to the inherent non-determinism and non-reproducibility of the underlying models (Hassan et al., 2024). Recent studies show that such non-determinism can persist even with fixed low temperature and Top-P settings in practical deployments (Atıl et al., 2025). Moreover, in many agentic contexts (e.g., creative assistants), this variability is expected and supports natural behavior, rather than indicating a defect (Krakowski, 2025; Hexmoor et al., 2026). While testing methodologies are well studied in traditional software engineering (Zhu et al., 2025; Zamprogno et al., 2022) and machine learning systems (Openja et al., 2024), the FM-based agent ecosystem remains underexplored. This study addresses this gap by presenting the first large-scale empirical investigation of unit testing practices in open-source agent frameworks and applications.
@@
Motivation: Agent ecosystem is a complex ecosystem of interacting components, from memory and tools to planning and coordination artifacts (Hettiarachchi, 2025). Understanding how testing effort is distributed across this heterogeneous architecture is vital for assessing whether critical components receive sufficient attention and for supporting the development of more systematic and effective testing practices. To create a durable analysis that transcends rapid evolution of agent frameworks and agentic applications, we opt to map observed components to a stable, canonical architectural model, e.g., JaCaMo framework (Boissier et al., 2020), which will allow us to assess whether testing effort is distributed effectively across fundamental roles. Then we investigate the relationship between the testing patterns identified in RQ1 and the underlying components of agent frameworks and applications.
@@ 2 A Motivational Example
Alex, a senior developer with a background in traditional software engineering, recently joined an ambitious AI startup. His first assignment is to lead the development of “HomeBot”, an autonomous Smart Home Management Agent. The agent’s core mandate is to manage IoT devices (e.g., thermostats, locks, and lights) and assist users with technical troubleshooting via natural language. To achieve this, Alex builds the system on a popular open-source agent framework that orchestrates a state-of-the-art Foundation Model (FM).
Phase 1: It works. In the initial phase, using a popular open-source agent framework, Alex quickly builds a prototype. The framework provides seamless integration with the FM. For example, when any user says “It’s too dark to read.”, the agent consistently interprets this intent and executes the turn_on_lights(brightness=100) tool. The initial results are impressive, and the team moves forward with deployment, confident in the power of their new AI-driven architecture.
Phase 2: The Unseen Decay. Weeks after launch, user feedback reveals a degradation in quality. Agents are becoming chatty, requiring increasing number of instructions to accomplish a task. Alex investigates and discovers a baffling pattern: for the exact same inputs, e.g., “It’s too dark to read”, the agent now exhibits non-deterministic behavior.
@@
    Behavior A (Desired): The agent silently executes the tool to turn on the lights.
@@
    Behavior B (Drift): The agent responds conversationally, “I can help with that. Would you like me to turn on the reading lamp?”
Both behaviors are plausible and expected from a conversational standpoint, yet they represent materially different system actions. Alex investigates and finds the root cause: a silent, unannounced update to the underlying FM has altered its interpretation of their prompts. The application is performing, yet the user-facing quality has declined. Alex realizes he has no automated regression suite capable of detecting this semantic drift; his standard unit tests expect deterministic string outputs, while the agent’s real-world behavior is intentionally non-deterministic.
Phase 3: The Component Explosion. Things become more complicated with the product team introducing two new features:
@@
    Contextual Troubleshooting (RAG): Users can ask, “What does the red blinking light on my heater mean?” This requires integrating a VectorDB to perform Retrieval-Augmented Generation (RAG) over technical manuals (Lewis et al., 2020).
@@
    Real-World Integration (MCP): Users can ask, “How much is my current electricity bill?” This requires the agent to interface with a third-party utility API via a Model Context Protocol (MCP) server (Hasan et al., 2025).
Alex’s testing challenges have now multiplied. Not only does he need to validate the non-deterministic output of the FM, but he must now also evaluate distinct components:
  • $\bullet$
    Testing RAG: How can he write a test to verify that the retrieved manual chunk is relevant to the blinking light error? A simple keyword match fails if the manual uses synonyms (e.g., flashing indicator).
  • $\bullet$
    Testing MCP: The utility API is a black box. If the agent fails to answer the billing question, is it because the FM failed to call the tool, or because the third-party MCP server changed its schema?
Phase 4: The Testing Gap. Drawing on his traditional software engineering background, Alex employs familiar techniques such as mocking and component-level testing. However, he quickly finds these practices difficult to apply in agentic systems where non-deterministic FM behavior is an expected design choice and components evolve independently. To seek guidance, he examines the testing suites of popular open-source agent frameworks and agentic applications. Instead of established practices, Alex finds a fragmented landscape of ad hoc scripts and coarse-grained end-to-end tests that offer little insight into component-level failures. There is no clear consensus on how practitioners test agentic systems, how they handle non-deterministic outputs, or which components receive focused testing. This gap motivates two fundamental questions:
@@
    How are testing efforts distributed across the components of AI agent frameworks and agentic applications? (explored in RQ2)
@@ 3.1 Foundations of Software Testing
Testing has been an integral part of software quality assurance. Software testing is typically defined as a process that systematically executes a program or system with the explicit objective of identifying and isolating defects (Myers, 2006). The focal point of this process is the Subject Under Test (SUT), which represents the specific software artifact being validated, ranging from atomic units like functions or classes to complex modules and entire services (Meszaros, 2007). These SUTs are exercised by test functions, also known as test cases, which serve as the atomic units of verification. A test function acts as an executable procedure that initializes a specific state or fixture, executes a command on the SUT, and verifies the outcome against specified requirements (Van Rompaey and Demeyer, 2008; Tao, 2009).
To structure these test functions effectively, developers often employ the Arrange-Act-Assert (AAA) pattern. This convention divides a test into three distinct phases: arranging the preconditions and inputs, acting on the SUT, and asserting that the resulting state matches expectations (Wei et al., 2022a). Within this structure, testing patterns offer recurring solutions to common design challenges. For example, Test Doubles (such as mocks and stubs) are frequently used during the arrangement phase to isolate the SUT from external dependencies, ensuring that tests remain deterministic and focused solely on the logic under verification (Zhu et al., 2025; Meszaros, 2007).
@@ 3.2 The Agent Ecosystem
An agent is an autonomous software entity that perceives its environment and performs actions to achieve specified objectives (Liu et al., 2023a). Classical agent models, such as the Belief-Desire-Intention (BDI) framework, conceptualize agent behavior as a reasoning cycle over internal states, goals, and plans (Rao et al., 1995). Contemporary agent implementations extend this paradigm by incorporating foundation models as the primary reasoning mechanism, supported by memory systems, planning logic, and interfaces for interacting with external tools and services (Park et al., 2023; Hettiarachchi, 2025). This integration enables agents to address open-ended tasks while introducing new forms of statefulness, non-determinism, and environmental interaction.
From a software engineering standpoint, the agent ecosystem can be conceptualized as having two distinct layers: (1) agent frameworks and (2) agentic applications. Agent frameworks provide reusable infrastructure for defining agents, managing state and memory, orchestrating tool usage, and coordinating interactions among agents (Wu et al., 2024a; Li et al., 2023). Agentic applications are concrete software systems built on top of these frameworks, where domain-specific workflows, prompts, and tool compositions are implemented. This distinction is important for testing because frameworks expose generic execution primitives and abstractions, whereas applications encode task-specific logic and integration behavior. Consequently, the nature and granularity of unit tests differ across these layers.
@@ 3.3 Canonical Agent Architecture
To enable consistent analysis across heterogeneous agent frameworks and applications, we adopt a canonical agent architecture as an abstract reference model. This architecture is grounded in established multi-agent system models that decompose systems into agents, environments, and organizational structures (Boissier et al., 2020). The Agent dimension captures internal state and control constructs, such as beliefs, goals, and plans. The Environment dimension encompasses the artifacts, resources, and observable signals that agents interact with to effect change. The Organization dimension defines the coordination structures, roles, and communication mechanisms that govern interactions among autonomous entities.
This classical model is further extended to account for architectural elements that have become central in modern FM-based ecosystems, particularly Registries. These components facilitate the dynamic discovery and invocation of tools or agents across system boundaries, bridging the gap between static definitions and runtime capabilities (Liu et al., 2025; Hasan et al., 2025). By serving as a unifying vocabulary, this canonical architecture allows for the identification of Subjects Under Test and the categorization of test targets independent of framework-specific abstractions. We provide the formal definitions of these architectural components and their detailed mappings to contemporary agent systems in Appendix B.